Two charges Q1 of 3 C and Q2 of 5 C are located at ri a) Find the vector force F12 on Qs due to Q2 in terms of unit vectors I, j, and k b) Extract the x, y and z components of force c) Find the unit vector in the direction of the force F12. d) Find the direction angles 21 + 4 j + 3k and r,-31+5-6k .
Answer #1

a)

Fi = force on charge Q1 by charge Q2 = k Q1 Q2 (\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} ) /(|\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} |)3

\underset{r_{1}}{\rightarrow} = 2 i + 4 j + 3 k

\underset{r_{2}}{\rightarrow}= 3 i + 5 j - 6 k

\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} = (2 i + 4 j + 3 k) - (3 i + 5 j - 6 k ) = - i - j + 9 k

|\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} | = sqrt((-1)2 + (-1)2 + (9)2) = 9.11

Fi = k Q1 Q2 (\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} ) /(|\underset{r_{1}}{\rightarrow}- \underset{r_{2}}{\rightarrow} |)3

Fi = (9 x 109) (3) (5) (- i - j + 9 k ) /(9.11)3

Fi = (1.8 x 108) (- i - j + 9 k )

Fi = - (1.8 x 108) i - (1.8 x 108) j + (16.2 x 108) k

b)

x- component : - (1.8 x 108)

y-component : - (1.8 x 108)

z-component : (16.2 x 108)

c)

|Fi| = sqrt((- (1.8 x 108) )2 + (- (1.8 x 108) )2 + (16.2 x 108)2) = 1.6 x 109

unit vector is given as

unit vector = Fi / |Fi| = (- (1.8 x 108) i - (1.8 x 108) j + (16.2 x 108) k) / (1.6 x 109 )